1.     Suppose that business travelers and vacationers have the following demand for airline tickets from New York to Boston:


As the price of tickets rises from $200 to $250, what is the price elasticity of demand for (i) business travelers and (ii) vacationers? 


1
Expert's answer
2021-10-14T16:40:54-0400



"PE_d =\frac{\frac{Q_2-Q_1}{Q_2+Q_1}/2}{\frac{P_2-P_1}{P_2+P_1}/2}"

(i) For business travellers:

"PE_d= \frac{\frac{2000-2100}{2000+2100}/2}{\frac{200-150}{200+150}/2}"

"PE_d= -0.17"

-0.17 is less than 1, demand is inelastic.


(ii)For vacationers:

"PE_d=\frac{\frac{800-1000}{800+1000}/2}{\frac{200-150}{200+150}/2}"

"PE_d=-0.78"

-0.78 is less than 1, demand is inelastic.